# On discount functions for economic model predictive control without terminal conditions

### v1

@@ Appendix A Technical Lemmas
and define $\delta_{9}(N)=2\bar{\lambda}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)$ and note that $\delta_{9}\in\mathcal{L}$ as $\beta(1)=0$ and $\beta$ is continuous. Now let us turn to $\Sigma_{6}$ and use Lemma A.3 to obtain that there exists $\delta_{10}\in\mathcal{L}$ with $\Sigma_{6}\leq\delta_{10}(N)$, where the assumptions are satisfied since $\lambda$ is continuous and the weak turnpike property holds not only for $x_{y}^{\star},u_{y}^{\star}$ but also for $\tilde{x}_{x}^{\star},\tilde{u}_{x}^{\star}$. The fact that the sum in $\Sigma_{6}$ goes only until $P-1$ and not until $N-1$ just implies that we get another $\frac{L_{\beta}}{N}(\sqrt[4]{N^{3}}+1)2\bar{\lambda}\in\mathcal{L}$ on top of $\delta_{10}$, as $N-P\leq\sqrt[4]{N^{3}}+1$. Next, we consider the part $\Sigma_{5}$ and construct a suboptimal candidate solution $\bar{u}\in\mathbb{U}^{N}(x)$ for $V_{N}^{\beta}(x)$ that follows the optimal solution $\tilde{u}_{x}^{\star}$ of $\tilde{V}_{N}^{\beta}(x)$ for $P$ steps, then follows the local controllability input of Assumption 7 that brings the state in $K\in\mathbb{I}_{[M_{1},M_{1}+p^{\star}-1]}$ steps [^1] Since $\tilde{x}_{x}^{\star}(P)$ and $x_{y}^{\star}(P)$ are both in the $\sigma_{2}(N)$ neighborhood of ${\Pi^{\star}_{\mathbb{X}}}$, but not necessarily in the $\sigma(N)$-neighborhood of the same ${\Pi^{\star}_{\mathbb{X}}}(j)$, we need at most $p^{\star}-1$ additional steps to have the same phase. to $x_{u_{y}^{\star}}(P,y)$, and then stays on the trajectory of $V_{N}^{\beta}(y)$ by applying $u_{y}^{\star}$. Hence,
$\Sigma_{5}$ $=\sum_{k=0}^{P-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)-J_{N}^{\beta}(x,\bar{u})+\sum_{k=P}^{P+K-1}{\textstyle\beta\big{(}\frac{k}{N}\big{)}}\ell\left(x_{\bar{u}}(k,x),\bar{u}(k)\right)$
$\qquad+\sum_{k=P}^{N-K-1}{\textstyle\beta\big{(}\frac{k+K}{N}\big{)}}\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)$
$\leq V_{N}^{\beta}(y)-V_{N}^{\beta}(x)+{\textstyle\beta\big{(}\frac{P}{N}\big{)}}K\bar{\ell},$
@@
where the last inequality holds as $\ell$ and $\beta$ are non-negative and $\beta$ is non-increasing. As $P\geq N-\sqrt[4]{N^{3}}-1$, we know that ${\textstyle\beta\big{(}\frac{P}{N}\big{)}}K\bar{\ell}\leq{\textstyle\beta\big{(}\frac{N-\sqrt[4]{N^{3}}-1}{N}\big{)}}K\bar{\ell}=:\delta_{12}(N)$ where $\delta_{12}\in\mathcal{L}$ as $\beta$ is continuous and $\beta(1)=0$. Finally, define $\delta_{3}(N):=\delta_{9}(N)+\delta_{9}(N)+\delta_{10}(N)+\delta_{12}(N)$ for $N\geq N_{0}$ and plug the bounds on $\Sigma_{5}$ and $\Sigma_{6}$ into (21) to obtain the desired result for $N\geq N_{0}$. As (18) holds with $\delta_{3}(N):=\max\{\delta_{3}(N_{0}),N_{0}(2\bar{\ell}+4\bar{\lambda})+2\bar{\lambda}\}$ for $N<N_{0}$, which guarantees $\delta_{3}\in\mathcal{L}$, the desired result follows for all $N\in\mathbb{N}$.

### v2

@@ 5 Conclusion and outlook
and define $\delta_{9}(N)=2\bar{\lambda}\beta\left({\textstyle 1-\frac{1}{\sqrt[4]{N}}-\frac{2}{N}}\right)$ and note that $\delta_{9}\in\mathcal{L}$ as $\beta(1)=0$ and $\beta$ is continuous. Now let us turn to $\Sigma_{6}$ and use Lemma A.3 to obtain that there exists $\delta_{10}\in\mathcal{L}$ with $\Sigma_{6}\leq\delta_{10}(N)$, where the assumptions are satisfied since $\lambda$ is continuous and the weak turnpike property holds not only for $x_{y}^{\star},u_{y}^{\star}$ but also for $\tilde{x}_{x}^{\star},\tilde{u}_{x}^{\star}$. The fact that the sum in $\Sigma_{6}$ goes only until $P-1$ and not until $N-1$ just implies that we get another $\frac{L_{\beta}}{N}(\sqrt[4]{N^{3}}+1)2\bar{\lambda}\in\mathcal{L}$ on top of $\delta_{10}$, as $N-P\leq\sqrt[4]{N^{3}}+1$. Next, we consider the part $\Sigma_{5}$ and construct a candidate solution $\bar{u}\in\mathbb{U}^{N}(x)$ for $V_{N}^{\beta}(x)$ which consists of the first $P$ steps of the optimal solution $\tilde{u}_{x}^{\star}$ of $\tilde{V}_{N}^{\beta}(x)$ and afterwards uses the local controllability to go to $x_{u_{y}^{\star}}(P,y)$ in $K\in\mathbb{I}_{[M_{1},M_{1}+p^{\star}-1]}$ steps [^1] Since $\tilde{x}_{x}^{\star}(P)$ and $x_{y}^{\star}(P)$ are both in the $\sigma_{2}(N)$ neighborhood of ${\Pi^{\star}_{\mathbb{X}}}$, but not necessarily in the $\sigma(N)$-neighborhood of the same ${\Pi^{\star}_{\mathbb{X}}}(j)$, we need at most $p^{\star}-1$ additional steps to have the same phase. and then applies $u_{y}^{\star}(k)$ starting at $k=P$ to remain on $V_{N}^{\beta}(y)$. Hence,
$\Sigma_{5}$ $=\sum_{k=0}^{P-1}{\textstyle\beta\left(\frac{k}{N}\right)}\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)-J_{N}^{\beta}(x,\bar{u})+\sum_{k=P}^{P+K-1}{\textstyle\beta\left(\frac{k}{N}\right)}\ell\left(x_{\bar{u}}(k,x),\bar{u}(k)\right)$
$\qquad+\sum_{k=P}^{N-K-1}{\textstyle\beta\left(\frac{k+K}{N}\right)}\ell\left(x_{y}^{\star}(k),u_{y}^{\star}(k)\right)$
$\leq V_{N}^{\beta}(y)-V_{N}^{\beta}(x)+{\textstyle\beta\left(\frac{P}{N}\right)}K\bar{\ell},$
@@
where the last inequality holds as $\ell$ and $\beta$ are non-negative and $\beta$ is non-increasing. As $P\geq N-\sqrt[4]{N^{3}}-1$, we know that ${\textstyle\beta\left(\frac{P}{N}\right)}K\bar{\ell}\leq{\textstyle\beta\left(\frac{N-\sqrt[4]{N^{3}}-1}{N}\right)}K\bar{\ell}=:\delta_{12}(N)$ where $\delta_{12}\in\mathcal{L}$ as $\beta$ is continuous and $\beta(1)=0$. Finally, define $\delta_{3}(N):=\delta_{9}(N)+\delta_{9}(N)+\delta_{10}(N)+\delta_{12}(N)$ for $N\geq N_{0}$ and plug the bounds on $\Sigma_{5}$ and $\Sigma_{6}$ into (19) to obtain the desired result for $N\geq N_{0}$. As (10) holds with $\delta_{3}(N):=\max\{\delta_{3}(N_{0}),N_{0}(2\bar{\ell}+4\bar{\lambda})+2\bar{\lambda}\}$ for $N<N_{0}$, which guarantees $\delta_{3}\in\mathcal{L}$, the desired result follows for all $N\in\mathbb{N}$. ∎